Fermi National Laboratory
 

Accelerator Update for February 21, 2007 - February 23, 2007

 

Wednesday February 21

The day shift began with Operations monitoring Tevatron (TeV) store 5238 (74.07E30) and the Recycler stash (175.2E10), with the Antiproton Source (Pbar) stacking (38.3E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

At 8:46 AM, a ComEd power glitch hit Fermi and caused the TeV to quench at sector C1, the loss of store 5238, and caused many devices to trip off. The quench was only a half-cell at C1. Operations began recovery.

Operations resumed sending beam to Pbar and NuMI at 9:03 AM.

Operations put the TeV through a dry squeeze at 11:31 AM.

Operations reported at 1:19 PM that cryo system technicians began work on a wet engine flywheel at TeV sector E4.

At 1:30 PM, Operations reported that a feeder 32 trip caused another power glitch. This glitch caused the CHL nitrogen plant to trip off and the loss of power to a portion of the Neutrino line. Later it was discovered that this also caused problems in Meson. The High Voltage Coordinator responded.

At 2:09 PM, an A1 transfer line vacuum pump tripped off and took down the MI beam permit. The pump trip also caused a TeV F0 vacuum burst.

Cryo System techs reported that they had completed their repairs on the TeV sector E4 wet engine.

Operations put the TeV into a dry squeeze at 3:18 PM.

At 4 PM, the TeV beam permit tripped due to a bad beam position in sector F1, but there was no beam in the TeV at the time. Experts investigated.

Operations put the TeV into shot setup at 4:11 PM.

Operations reported at 4:13 PM that the Recycler had stashed antiprotons twice during the day shift.

TeV experts began a 30-minute sextupole study at 4:33 PM.

Operations established store 5240 at 6:36 PM with an initial average luminosity of 221.39E30.

At 6:48 PM, a Booster RF station (BRF6) tripped off for the first of two times in 30 minutes.

Operations reported at 11:54 PM that the Recycler had stashed twice during the evening shift.

Thursday February 22

The midnight shift began with Operations monitoring store 5240 (102.87E30) and the stash (121.0E10), with Pbar stacking (16.0E10), and with NuMI taking beam.

Operators search and secured the Meson MTest area at 1:19 AM in preparation for the new MTest experiment (T953) that will begin taking beam later today. They will run from Thursday through Sunday.

At 3:35 AM, Operations reported a problem with a Meson CAMAC crate that was overheating. Operators replaced its power supply.

At 5:05 AM, BRF6 tripped off. It reset.

At 5:08 AM, a focusing quadrupole magnet’s (E:QF112) current drifted below its lower limit and tripped off the NuMI beam permit. Everything reset okay.

At 5:21 AM, a defocusing quadrupole’s (E:QD108) current drifted above its upper limit and tripped off the NuMI beam permit. Everything reset okay.

At 5:22 AM, a Linac RF station (LRF5) tripped off for the first of a few times over an hour and 20 minutes. Turning the station on and back to its normal gradient each time was a slow process.

At 5:38 AM, Operations reported that the Meson MW1W power supply (the MTest critical device) had tripped off. Operations called in a power supply expert.

At 6:43 AM, LRF5 tripped off. Operators managed to restore the station after a long turn on ritual.

The MTest experiment reported at 6:58 AM that they were ready for beam. MW1W work continued, which was holding off their beam.

At 7:32 AM, Operations reported that the Meson vacuum was not good. An expert suspected that the feeder 32 trip had caused the problem. Operators prepared the area for access to investigate.

Operations reported at 8:07 AM that the Recycler had stashed twice during the midnight shift.

At 9:48 AM, BRF6 tripped off. It reset.

Linac experts reported at 10:06 AM that they adjusted the high voltage and current on LRF5. They also reported that the station’s power amplifier tube was weak and eventually would have to be replaced.

Vacuum techs reported at 11:13 AM that they found several vacuum pumps in Meson that had not reset from the power glitch on Thursday. They reset the pumps and fixed one that had failed.

Operations established beam to the Meson MTest experiment at 12:11 PM.

The MTest experiment (from the University of Iowa) intends to investigate various configurations of Cerenkov light generators and tubes. This work will lead to developing detectors that use Cerenkov light and phototubes as active elements.

What is Cerenkov light?

Pavel Alekseyevich Cerenkov received a Nobel Prize in 1958 for explaining the blue glow seen in nuclear reactors. The speed of light is a constant in a vacuum and cannot be exceeded. However, the speed of light in a material is much slower. When a particle (usually an electron) passes through a material faster than light can pass through it, it disrupts the material and causes photons to be emitted. If this disruption continues traveling faster than light in the material (picture a sonic boom) it causes a measurable radiation. The mass of a particle traveling at a known speed can be found by measuring the angle of the emitted Cerenkov radiation. This provides an important tool for particle detection.

Pbar experts began a reverse proton study at 1:38 PM.

At 3:07 PM, LRF5 tripped off for the first of three times in 30 minutes. It reset each time.

Operations reported at 4:15 PM that the Recycler had stashed twice during the day shift.

Operations reported at 4:22 PM that MI extracted to NuMI a record intensity of 4.04E13 with one pulse. This was accomplished with 11-batch slipstacking.

At 4:49 PM, Operations reported that NuMI beam resumed. It had been off for about 10 minutes due to vacuum problems.

Pbar studies ended at 5:41 PM.

A TeV expert began an end-of-store study at 7:04 PM.

Operations terminated store 5240 at 8 PM.

Operations put the TeV into shot setup at 8:15 PM.

The MTest experiment reported at 9:05 PM that they were done with beam for the night.

At 9:09 AM, Operations reported that the TeV lost its loaded protons due to a loss of communication with a crate ($C3) at sector C3. Operators investigated and found that the crate’s power supply had failed. Operators had the power supply replaced by 10:09 PM.

At 10:10 PM, LRF5 tripped off. Operators had it up and running within 11 minutes.

Operations put the TeV through a dry squeeze at 10:32 PM.

At 10:49 AM, the NuMI beam permit tripped off due to a bad beam position.

Operations put the TeV into another shot setup at 10:57 PM.

At 11:26 PM, a MI crate power supply ($15) failed. Operators had the supply replaced by 11:41 PM.

Friday February 23

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (62.7E10) and stack (96.0E10), with MTest off, and with NuMI taking beam.

Operations established store 5242 at 12:53 AM with an initial luminosity of 244.3E30.

At 1:47 AM, KRF5 tripped off due to a water flow indication. It reset.

At 2:11 AM, BRF6 tripped off. It reset.

At 4:03 AM, MIRF12 tripped off for the first of four times in an hour. It reset each time. The first and last trip caused beam quality problems with the Pbar Debuncher and NuMI.

At 6:14 AM, Operations reported that MIRF12 continued to trip off and be reset. At 6:32 AM, it tripped off and would not reset. Operators called a RF expert. Later, operators bypassed the station.

Operations established beam to MTest at 6:36 AM.

At 6:40 AM, KRF1 tripped off on a LCW flow indications. It reset. A few minutes later the station’s RF wandered out of phase. It came back a minute later.

Operations reported at 8:17 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Friday and the Weekend

The plans for today are to stack, stash, and establish a new store this evening. The Run Coordinator will allow Linac experts to replace the PA for LRF5 and Booster RF experts to repair BRF14 this afternoon. He will allow TeV experts to conduct a crystal collimator end-of-store study on Sunday.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/26/2007    email Fermilab